OpenStudy (anonymous):

simplify: log[7](3)+ 3log[7]2

Parth (parthkohli):

\[\log_7 3 + 3\log_7 2 \iff \log_7 3 + \log_7 2^3 \iff \log_7(3\times 8)\cdots\]

Parth (parthkohli):

Use the properties:\[n\log a = \log a^n\]\[\log a +\log b = \log(ab)\]